a magic rectangular box shrinks ½ length and 1/3 width when its owner wises something.after three such wishes the suface area was 4cm2.what would be the original length if the original width was 9cm.

Grade:9

1 Answers

Nishant Vora IIT Patna
9 years ago
Let original Length = L cm
Original width = 9cm

After one wise
Length = L- L/2 = L/2
width= 9 – 9/3 = 6

After 2nd wise,
Length = L/2 – L/4 = L/4
width= 6 – 6/3 = 4

After 3rd wise
Length = L/4 – L/16 =3L/16
width= 4 – 4/3 = 8/3

Area = 3L/16 * 8/3 = L/2 = 4cm2

So L = 8 cm